# QTIP: Quantization with Trellises and Incoherence Processing

### v1

@@ Abstract
Post-training quantization (PTQ) reduces the memory footprint of LLMs by quantizing weights to low-precision datatypes. Since LLM inference is usually memory-bound, PTQ methods can improve inference throughput. Recent state-of-the-art PTQ approaches have converged on using vector quantization (VQ) to quantize multiple weights at once, which improves information utilization through better shaping. However, VQ requires a codebook with size exponential in the dimension. This limits current VQ-based PTQ works to low VQ dimensions ($\leq 8$) that in turn limit quantization quality. Here, we introduce QTIP, which instead uses trellis coded quantization (TCQ) to achieve ultra-high-dimensional quantization. TCQ uses a stateful decoder that separates the codebook size from the bitrate and effective dimension. QTIP introduces a spectrum of lookup-only to computed lookup-free trellis codes designed for a hardware-efficient “bitshift” trellis structure; these codes achieve state-of-the-art results in both quantization quality and inference speed.
@@ 1 Introduction
Large language models (LLMs) have accelerated advancements in fields ranging from natural language processing [34] to scientific modeling [28]. However, the largest LLMs have hundreds of billions of parameters that can take over a terabyte of memory to load in half-precision; this size poses significant challenges for the practical deployment of LLMs [33, 18, 2]. For example, small-batch autoregressive decoding, a common form of inference for LLMs, is memory bound [35]. Even on a modern datacenter GPU with $\approx 3$TB/s memory bandwidth, a large LLM ($\gtrapprox 200$GB) can only be directly run at $\lessapprox 20$ tokens per second and requires multiple devices [4]. One way to accelerate inference is by compressing LLMs. This directly reduces the memory footprint of the model and increases the theoretical maximum inference throughput on any given machine.
One form of compression, weight-only post-training quantization (PTQ), quantizes trained model weights to lower precision datatypes [10, 35, 6]. The latest state-of-the-art weight-only PTQ methods, QuIP##\## and AQLM, use vector quantization (VQ) to achieve high-quality 2-bit models [35, 12]. In VQ, a vector $x\in\mathbb{R}^{d}$ is quantized to one of $2^{kd}$ vectors in $\mathbb{R}^{d}$ that form a codebook $C\in\mathbb{R}^{2^{kd}\times d}$. A higher vector dimension $d$ allows for better codebook shaping and packing density, improving information utilization [20]. However, VQ requires exponential time and space in both the bitrate and dimension, limiting its practicality. During quantization, VQ costs $O(2^{kd}d)$ time to perform nearest-neighbor rounding to $C$, and during inference, $C$ must fit in hardware cache for fast lookups. This exponential scaling limits how high $d$ can be and thus the advantages of VQ over scalar quantization.
To address this limitation, we propose QTIP, which uses trellis-coded quantization (TCQ) to enable tractable ultra-high-dimensional ($>100$) quantization and improves quantization quality over prior VQ-based approaches. In the simplest scalar form of TCQ, a length-$T$ sequence $S$ is statefully quantized using a trellis – a directed graph $G$ with $2^{L}$ nodes, each with $2^{k}$ incoming and outgoing edges and a scalar value [24]. The reconstructed sequence $\hat{S}$ corresponds to the node values of a length-$T$ walk on $G$, and quantization finds the walk that minimizes some distortion metric on $S$ and $\hat{S}$. Since neighboring entries in $\hat{S}$ are connected by one of $2^{k}$ edges, we only need to store which edge an entry came from, which takes $k$ bits. For additive distortion metrics such as squared error, the optimal $\hat{S}$ can be found with the Viterbi algorithm, which runs in $O(2^{L}T)$ time [14, 24]. This means that the cost of quantization is independent of the bitrate $k$ and linear in the sequence dimension $T$, enabling tractable high dimensional quantization.
@@
    Simultaneously achieves state-of-the-art weight-only LLM PTQ quality and fast inference through a combination of hardware-efficient trellis and codebook design.
@@
    Introduces multiple novel hardware-efficient ($\leq 4$ instructions per weight) computed and hybrid lookup-computed random Gaussian codes for TCQ on i.i.d. Gaussian sources.
@@
Figure 1: QTIP performs ultra-high dimensional ($>100$) quantization by using Trellis Coded Quantization, which has linear cost in dimension. This enables QTIP to outperform Vector Quantization-based approaches (QuIP##\##, AQLM) that are limited to small dimensions.
@@ 3.1.1 Lookup-Free Computed Codes
Algorithm 1 (1MAD) first runs a linear congruential generator (LCG) to produce a pseudorandom 32-bit word. This requires 2 instructions (MAD and &). It then sums the 32-bit word as four 8-bit unsigned integers; this sum is approximately Gaussian distributed. This requires 1 instruction (vabsdiff4). Finally, this sum must be scaled and shifted (another MAD). Although there are only $2^{10}$ representable values even when $L>10$, this does not empirically affect quantization quality. 1MAD requires choosing $a$ and $b$ to avoid strong correlations; we set $a=34038481$ and $b=76625530$ (Figure 3 LC).
@@ 4 Experiments
Here, we present experiments quantizing the Llama family of models with QTIP [33, 34, 26]. These models offer strong performance across a wide range of sizes, allowing us to compare how different quantization methods perform and scale. We primarily compare QTIP against QuIP##\## and AQLM. For Llama 1, we include GPTVQ-2D instead of AQLM since AQLM does not publish Llama 1 numbers [36]. GPTVQ-2D performs 2D VQ inside GPTQ and offers strong performance. These methods outperform scalar quantization methods including GPTQ, AWQ, and OmniQuant; comparisons to those methods can be found in QuIP##\## and AQLM [21, 15, 30, 35, 12]. We mainly focus on the hybrid code (Section 4.2) since it is tailored for modern GPUs, and present a full suite of results for it. For the computed codes (Section 4.1), we present results for Llama 2.
@@ 4.1 Lookup-Free Computed Codes
Table 3: Wikitext2 and C4 perplexity ($\downarrow$), ctx. 4096, QTIP with pure-computed codes. Even without fine-tuning, pure-computed QTIP outperforms QuIP##\## with fine-tuning at almost all models and sizes.
@@ 4.2 Hybrid Lookup-Computed Codes
FP16 16 51.1 77.7 81.1 77 16 45.6 73.3 73.5 69.6 16 40 69.3 78.5 67.3
AQLM 4.07 51.0 78.1 81.4 76.9 3.94 43.9 72.2 78.6 70.4 4.04 40.3 68.9 77.7 67.3
QuIP# 4 50.6 78.1 81.4 77.1 4.00 45.5 73.9 78.9 69.9 4.00 40.5 69.1 78.4 67.6
QTIP 4 50.0 77.6 81.5 77.0 4.00 45.9 73.2 78.6 69.9 4.00 40.4 69.2 78.5 67.4
AQLM 3.01 50.0 77.6 81.3 77.2 3.03 43.6 73.5 77.8 67.6 3.04 38.7 67.8 76.6 68.4
QuIP# 3 50.0 77.7 81.4 76.4 3.00 44.0 72.5 78.4 69.1 3.00 39.2 68.4 77.3 66.5
QTIP 3 49.8 77.5 81.2 76.3 3.00 43.6 72.5 78.2 69.5 3.00 39.8 69.7 78.0 66.8
AQLM 2.07 47.9 77.7 80.4 75.9 1.97 38.5 67.0 75.1 69.5 2.02 33.6 62.8 73.5 64.6
QuIP# 2 48.7 77.3 80.3 75.9 2.00 39.5 69.3 77.3 67.7 2.00 34.6 64.6 75.1 64.9
QTIP 2 48.1 76.9 80.1 76.5 2.00 39.2 70.6 77.8 71.0 2.00 35.3 63.9 75.3 66.7
@@
Table 5 shows zeroshot results computed with LM Eval, which are slightly random; QTIP generally matches or exceeds QuIP##\## and AQLM on these tasks [16]. Table 6 contains results on Llama 3. Like other works, we have observed that Hessian-based adaptive rounding (GPTQ/LDLQ/BlockLDLQ) performs worse than expected on Llama 3 [17]. Since the contribution and focus of this work is what to round with (TCQ) and not how to round (BlockLDLQ), we only compare against the proximal baseline QuIP##\##, which uses BlockLDLQ with VQ. QTIP significantly improves upon QuIP##\## at all model sizes and bitrates, once again showing the dimensionality advantage of TCQ over VQ.
@@ 4.3 Inference Speed
Table 7: RTX4090 (1TB/s) Throughput
2-7B 2-70B
Mthd. tok/s
matvec
mem bw
tok/s
matvec
mem bw
FP16 33.1 OOM OOM
AQLM 20.6 8.27
QuIP# 106.3 697 GB/s 25.9 867 GB/s
QTIP 105.2 628 GB/s 25.8 840 GB/s
Table 7 shows the batch size 1 inference speed of QTIP, QuIP##\##, and AQLM on Llama 2 7b and 70b without matrix fusion. Here, the design choices of QTIP and QuIP##\## become apparent. Whereas AQLM uses a codebook that is too large to fit in cache and thus prevents fast inference, both QTIP and QuIP##\## achieve significant speedups over FP16. According to NVIDIA Nsight Systems, the matrix-vector multiply kernels of QuIP##\## and QTIP both run at near maximum memory bandwidth [29]. This means that a significant part of the gap between their kernel memory bandwidths and effective tok/s is due to poor kernel fusion in the rest of the inference pipeline, which could easily be optimized to further improve throughput. Finally, while it is impressive that both QuIP##\## and QTIP are 3-5$\times$ faster than AQLM, it is even more impressive that QTIP is able to match QuIP##\##’s throughput with an effective dimension size of 256, or 32×32$\times$32 × larger than QuIP##\##’s. This means that the improved quantization quality of QTIP comes with no additional inference-time cost.
@@ A.1 Additional Results
A.1.1 Lookup-Only Codes

### v2

@@ Abstract
Post-training quantization (PTQ) reduces the memory footprint of LLMs by quantizing weights to low-precision datatypes. Since LLM inference is usually memory-bound, PTQ methods can improve inference throughput. Recent state-of-the-art PTQ approaches use vector quantization (VQ) to quantize multiple weights at once, which improves information utilization through better shaping. However, VQ requires a codebook with size exponential in the dimension. This limits current VQ-based PTQ works to low VQ dimensions ($\leq 8$) that in turn limit quantization quality. Here, we introduce QTIP, which instead uses trellis coded quantization (TCQ) to achieve ultra-high-dimensional quantization. TCQ uses a stateful decoder that separates the codebook size from the bitrate and effective dimension. QTIP introduces a spectrum of lookup-only to computed lookup-free trellis codes designed for a hardware-efficient “bitshift” trellis structure; these codes achieve state-of-the-art results in both quantization quality and inference speed.
@@ 1 Introduction
Large language models (LLMs) have accelerated advancements in fields ranging from natural language processing [34] to scientific modeling [28]. However, the largest LLMs have hundreds of billions of parameters that can take over a terabyte of memory to load in half-precision; this size poses significant challenges for the practical deployment of LLMs [33, 17, 2]. For example, small-batch autoregressive decoding, a common form of inference for LLMs, is memory bound [35]. Even on a modern datacenter GPU with $\approx 3$TB/s memory bandwidth, a large LLM ($>200$GB) can only be directly run at $<20$ tokens per second and may require multiple devices [35]. One way to accelerate inference is by compressing LLMs. This directly reduces the memory footprint of the model and increases the theoretical maximum inference throughput on any given machine.
One form of compression, weight-only post-training quantization (PTQ), quantizes trained model weights to lower precision datatypes [9, 35, 5]. The latest state-of-the-art weight-only PTQ methods, QuIP##\## and AQLM, use vector quantization (VQ) to achieve high-quality 2-bit models [35, 11]. In VQ, a vector $x\in\mathbb{R}^{d}$ is quantized to one of $2^{kd}$ vectors in $\mathbb{R}^{d}$ that form a codebook $C\in\mathbb{R}^{2^{kd}\times d}$. A higher vector dimension $d$ allows for better codebook shaping and packing density, improving information utilization [19]. However, unstructured VQ requires exponential time and space in both the bitrate and dimension, limiting its practicality. During quantization, VQ costs $O(2^{kd}d)$ time to perform nearest-neighbor rounding to $C$, and during inference, $C$ must fit in hardware cache for fast lookups. This exponential scaling limits how high $d$ can be and thus the advantages of VQ over scalar quantization.
To address this limitation, we propose QTIP, which uses trellis-coded quantization (TCQ) to enable tractable ultra-high-dimensional ($>100$) quantization and improve quantization quality over prior VQ-based approaches. In the simplest scalar form of TCQ, a length-$T$ sequence $S$ is statefully quantized using a trellis – a directed graph $G$ with $2^{L}$ nodes, each with $2^{k}$ incoming and outgoing edges and a scalar value [24]. The reconstructed sequence $\hat{S}$ corresponds to the node values of a length-$T$ walk on $G$, and quantization finds the walk that minimizes some distortion metric on $S$ and $\hat{S}$. Since neighboring entries in $\hat{S}$ are connected by one of $2^{k}$ edges, we only need to store which edge an entry came from, which takes $k$ bits. For additive distortion metrics such as squared error, the optimal $\hat{S}$ can be found with the Viterbi algorithm, which runs in $O(2^{L}T)$ time [13, 24]. This means that the cost of quantization is independent of the bitrate $k$ and linear in the sequence dimension $T$, enabling tractable high dimensional quantization.
@@
    Achieves a state-of-the-art combination of weight-only LLM PTQ quality and fast inference through hardware-efficient trellis and codebook design.
@@
    Introduces multiple novel hardware-efficient ($\leq 4$ instructions per weight) compute-based random Gaussian codes for TCQ on i.i.d. Gaussian sources.
@@
Figure 1: QTIP performs ultra-high dimensional ($>100$) quantization by using Trellis Coded Quantization, which has linear cost in dimension. This enables QTIP to outperform Vector Quantization-based approaches (QuIP##\##, AQLM) that are limited to low dimensions. With QTIP, 2 bit models scale better than theoretically optimal 4 bit models.
@@ 3.1.1 Lookup-Free Computed Codes
Algorithm 1 (1MAD) first runs a linear congruential generator (LCG) to produce a pseudorandom 32-bit word [32]. This requires 2 instructions (MAD and &). It then sums the 32-bit word as four 8-bit unsigned integers; this sum is approximately Gaussian distributed. This requires 1 instruction (vabsdiff4). Finally, this sum must be scaled and shifted (another MAD). Although there are only $2^{10}$ representable values even when $L>10$, this does not empirically affect quantization quality. 1MAD requires choosing $a$ and $b$ to avoid strong correlations; we set $a=34038481$ and $b=76625530$ (Figure 3 LC).
@@ 4 Experiments
Here, we present experiments quantizing the Llama family of models with QTIP [33, 34, 26]. These models offer strong performance across a wide range of sizes, allowing us to compare how different quantization methods perform and scale. We primarily compare QTIP against QuIP##\## and AQLM. For Llama 1, we include GPTVQ-2D instead of AQLM since AQLM does not publish Llama 1 numbers [36]. GPTVQ-2D performs 2D VQ inside GPTQ and offers strong performance. These methods outperform scalar quantization methods including GPTQ, AWQ, and OmniQuant; comparisons to those methods can be found in QuIP##\## and AQLM [20, 14, 29, 35, 11]. We mainly focus on the hybrid code (Section 4) since it is tailored for modern GPUs, and present a full suite of results for it. For the computed codes (Section 4.1), we present results for Llama 2.
@@ 4.1 Lookup-Free Computed Codes
Table 3: Wikitext2 and C4 perplexity ($\downarrow$), ctx. 4096, QTIP with pure-computed codes. Even without fine-tuning, pure-computed QTIP outperforms QuIP##\## and AQLM, both of which use fine-tuning, at almost all models sizes.
@@ 4.2 Hybrid Lookup-Computed Codes
Table 4: Batch size 1 decoding throughput on a RTX6000 Ada (960GB/s mem. BW).
Method Bits 2-7B Tok/s 2-70B Tok/s
FP16 16 55.9 OOM
AQLM 2 81.5 8.78
QuIP# 2 186 22.2
QTIP 2 188 23.5
QTIP 3 161 19.1
QTIP 4 140 16.3
@@
FP16 16 51.1 77.7 81.1 77.0 16 45.6 73.3 73.5 69.6 16 40.0 69.3 78.5 67.3
AQLM 4.14 50.7 77.3 81.5 76.5 3.94 44.8 73.3 78.4 69.9 4.04 41.0 70.2 78.2 67.3
QuIP# 4 50.5 77.7 81.4 77.3 4 43.6 71.3 78.7 69.6 4 40.4 68.6 78.5 67.4
QTIP 4 50.0 77.8 81.3 76.9 4 44.8 73.6 78.9 69.9 4 40.0 68.9 78.4 67.1
AQLM 3.01 50.3 78.0 80.7 75.3 3.03 42.8 72.9 78.5 68.8 3.04 38.5 66.8 77.3 65.4
QuIP# 3 50.9 77.6 81.4 76.1 3 44.0 72.5 78.4 69.1 3 39.2 68.4 77.3 66.5
QTIP 3 50.3 78.2 80.6 77.0 3 44.0 72.8 78.0 69.5 3 38.9 68.1 78.1 66.9
AQLM 2.07 47.9 77.7 80.4 75.9 1.97 38.8 69.3 75.9 68.8 2.02 32.8 63.7 74.8 65.7
QuIP# 2 47.6 77.1 79.5 74.6 2 39.6 69.0 77.3 67.4 2 35.2 65.3 75.4 64.9
QTIP 2 48.0 76.3 80.2 75.1 2 41.4 70.8 77.3 67.6 2 35.7 65.6 75.9 64.7
@@
Table 8: Llama 3.1 instruct-tuned model results (ctx. 8192 for perplexity). QTIP performs well at all model sizes and generally outperforms PV-Tuning, a recent quantization method that focuses on fine-tuning algorithms. The zeroshot results in this table use LM Eval 0.4.4 and the “standard” versions of each task instead of the Meta versions in [26].
Ppl. ($\uparrow$) Zeroshot ($\downarrow$)
Bits W2 ArcC ArcE Hswag PiQA
Meta "FP8" 16 Attn. / 8 MLP 1.70 61.6 81.4 67.1 83.8
QTIP 4 1.79 61.3 80.9 66.7 84.2
QTIP 3 2.05 61.5 81.4 66.8 83.5
3.1 405B Inst. QTIP 2 3.29 60.7 81.1 65.4 82.2
BF16 16 3.52 56.7 75.6 61.5 82.8
QTIP 4 3.73 56.3 75.8 61.4 83.0
QTIP 3 4.12 55.1 75.1 60.8 82.6
QTIP 2 5.08 54.4 72.6 59.4 82.5
3.1 70B Inst. PV-Tuning 2.01 5.70 52.7 72.2 60.2 82.6
BF16 16 6.50 51.6 77.8 57.7 80.0
QTIP 4 6.61 50.7 78.0 57.5 80.1
QTIP 3 6.80 50.4 77.7 56.9 79.3
QTIP 2 7.82 45.1 75.6 54.5 79.0
3.1 8B Inst. PV-Tuning 2.07 8.45 46.2 75.4 54.4 78.7
Table 9: Llama 3.2 instruct-tuned results when quantizing to 4 bits (ctx. 8192 for perplexity). Even on extremely small models, QTIP is still able to achieve meaningful compression without sacrificing quality. This table uses the same LM Eval setup as Table 8.
Ppl ($\downarrow$) Zeroshot ($\uparrow$)
Size (GB) W2 ArcC ArcE Hswag PiQA
BF16 6 9.58 43.3 74.3 52.2 75.7
3B QTIP 2.1 9.77 43.5 74.3 51.9 75.1
BF16 2.4 11.57 36.0 68.5 45.2 74.2
1B QTIP 0.97 11.93 34.8 68.4 44.5 73.3
@@
Table 6 shows zeroshot results computed with LM Eval, which are slightly random; QTIP generally matches or exceeds QuIP##\## and AQLM on these tasks [15]. Table 7 contains results on Llama 3. Like other works, we have observed that Llama 3 (especially 70B base) is harder to quantize than Llama 2 [16]. Since the contribution and focus of this work is what to round with (TCQ) and not how to round (BlockLDLQ), we only compare against the proximal baseline QuIP##\##, which uses BlockLDLQ with VQ. QTIP significantly improves upon QuIP##\## at all model sizes and bitrates, once again showing the dimensionality advantage of TCQ over VQ. Table 8 shows results for Llama 3.1 instruct-tuned models, including Llama 3.1 405B. At all sizes, QTIP achieves strong results. Notably, QTIP is able to match or exceed PV-Tuning, a recent quantization method that focuses on better fine-tuning algorithms [22]. However, PV-Tuning is based off of AQLM and inherits its slow inference speed, making it significantly slower than QTIP. Finally, Table 9 shows results for quantizing Llama 3.2 instruct-tuned models to 4 bits. Since the embedding layers are very large relative to the decoder layers for small Llama 3 models ($\approx 500-750$MB), quantizing the decoder layers to fewer than 4 bits does not make a significant difference on the final model size. Here, QTIP is still able to achieve a meaningful end-to-end compression rate (2.5-3X) without degrading the final model.
@@ 4.3 Inference Speed
Table 4 shows the batch size 1 inference speed of QTIP, QuIP##\##, and AQLM on Llama 2 7B and 70B with matrix fusion. Here, the design choices of QTIP and QuIP##\## become apparent. Whereas AQLM uses a codebook that is too large to fit in cache and thus prevents fast inference, both QTIP and QuIP##\## achieve significant speedups over FP16. Furthermore, while it is impressive that both QuIP##\## and QTIP are >2×>2$\times$> 2 × faster than AQLM, it is even more impressive that QTIP is able to match QuIP##\##’s throughput with an effective dimension size of 256, or 32×32$\times$32 × larger than QuIP##\##’s. This means that the improved quantization quality of QTIP comes with no additional inference-time cost. Although our empirical throughput numbers were timed on NVIDIA GPUs, QTIP can be fast on a broad class of accelerators due to its flexibility. QTIP only requires generating a pseudorandom Gaussian efficiently, and can work on devices with no cache as well as devices with lookup hardware. For example, if we were using a ARMv8 CPU, we could use the vqtbl4q_u8 NEON intrinsic to look up 16 indices in a 64-entry codebook. This would let us use a 6 bit 1D codebook with the HYB code (Q=6, V=1). Quantizing Llama 2 7B to 2 bits with this setup and w/out fine-tuning gives 6.89 Wikitext2 perplexity – essentially the same state-of-the-art quality as 3INST.
@@ 5 Conclusion
Acknowledgements
C.D. was supported by NSF-2046760 CAREER. We thank Together AI for compute resources.
@@ A.1 Additional Results
A.1.1 Ablations on Trellis Size
Table 10 shows an ablation on $L$ for quantizing Llama 2 7B with $K=2,V=1$, the bitshift trellis, a pure-lookup codebook, and no fine-tuning. $L=8$ is the largest $L$ achievable if we had to store the trellis and codebook in the same amount of cache as the HYB code (2KiB). $L=10$ is the largest $L$ achievable if we only had to store the codebook. As expected, increasing $L$ improves quality. Table 10 also shows very little difference between an equal-sized LUT codebook and QTIP’s codes, meaning that QTIP isn’t sacrificing quality for speed. However, an equal-sized LUT would need >10×>10$\times$> 10 × more cache than the latest GPUs have, making the bitshift trellis and compute-based codes necessary to achieve both quality and speed. Table 11 shows an ablation on $V$ with $L=12$ and 16, $K=2$, and the same settings as Table 10. Increasing $V$ generally decreases quality, but this can be recovered with a larger $L$. It is hard to measure $V$’s impact on decoding speed since this is highly implementation and hardware dependent, so $V$ is more of a user-chosen hyperparameter.
Table 10: Ablation on $L$ when quantizing Llama 2 7B to 2 bits ($K=2$ and $V=1$).
L Trellis Size CB size total size W2 C4
QuIP# - 8Kb 8Kb 8.22 11.0
8 8.19 Kb 4.10 Kb 12.29 Kb 7.83 10.3
10 40.96 Kb 16.38 Kb 57.34 Kb 7.49 9.67
12 196.61 Kb 65.54 Kb 262.14 Kb 6.97 9.21
16 4.19 Mb 1.05 Mb 5.24 Mb 6.83 8.92
16 Bitshift 3INST 0Kb 6.82 8.96
Table 11: Ablation on $V$ when quantizing Llama 2 7B to 2 bits ($K=2$).
Codebook L V W2 C4
LUT 12 1 6.97 9.21
LUT 12 2 7.09 9.24
LUT 12 4 7.55 9.88
LUT 16 1 6.83 8.92
LUT 16 2 6.79 8.97
QTIP HYB (no FT) 16 2 6.83 8.97
LUT 16 4 6.92 9.07
A.1.2 Zeroshot Results
@@
A.1.3 Lookup-Only Codes
@@
A.1.4 Decoding Speed on Different GPUs
Table 17: Decoding speed on different Ampere and Lovelace GPUs.
GPU Model Model 2-bit tok/s 3-bit tok/s 4-bit tok/s FP16 tok/s
RTX 3090 2-7 127 119 109 52.5
RTX 3090 2-70 15.3 OOM OOM OOM
RTX A6000 Ampere 2-7 116 106 95 43.5
RTX A6000 Ampere 2-70 15.0 13.1 11.7 OOM
RTX 6000 Ada 2-7 188 161 140 55.9
RTX 6000 Ada 2-70 23.5 19.1 16.3 OOM